OpenStudy (anonymous):

A differentiable function f(x) is defined such that, for all values of x in its domain, (i will draw the equation). (a) what is the domain. (b) for what values of x is f(x)=3? (c)show that f'(x)=3^(2)f(x) (d)solve the differential equation in (c) to find f(x) in terms of x only.

OpenStudy (anonymous):

\[f(x)=3+\int\limits_{8}^{x^3}f(\sqrt[3]{t})dt\]

OpenStudy (anonymous):

\[f(x)=3\iff \int_8^{x^3}f(\sqrt[3]{t})dt=0\iff x^3=8\iff x=2\]